OpenStudy (anonymous):

the length of a rectangle is 5 centimeters more than the width. the area of the rectangle is 36 square centimeters. what is the length? (the width is x+5cm)

OpenStudy (anonymous):

Area = length x width (w+5)*w= 36 w^2+5w-36=0 (w-4)*(w+9)=0 w=4,w+5=9, length=9

OpenStudy (anonymous):

thanks so much! i had it backwards!

OpenStudy (anonymous):

Your welcome.